(1 point) Is the number of games won by a major league baseball team in a season related to the team batting average? The table below shows the number of games won and the batting average (in thousandths) of 8 teams. Team Games Won Batting Average 94 106 74 104 97 83 120 89 267 270 285 266 288 259 266 273 4 Using games won as the explanatory variable x, do the following: (a) The linear correlation coefficient is r (b) The equation of the least squares regression line is y
Answer #1

The statistical software output for this problem is:

Simple linear regression results:
Dependent Variable: Batting Average
Independent Variable: Games Won
Batting Average = 292.30825 - 0.21442765 Games Won
Sample size: 8
R (correlation coefficient) = -0.30927707
R-sq = 0.095652304
Estimate of error standard deviation: 10.238602

Parameter estimates:

Parameter Estimate Std. Err. Alternative DF T-Stat P-value
Intercept 292.30825 26.059229 ≠ 0 6 11.217072 <0.0001
Slope -0.21442765 0.26916907 ≠ 0 6 -0.7966281 0.456


Analysis of variance table for regression model:

Source DF SS MS F-stat P-value
Model 1 66.526177 66.526177 0.63461634 0.456
Error 6 628.97382 104.82897
Total 7 695.5

Hence,

a) Linear correlation coefficient = - 0.3093

b) Equation of regression line:

\widehat{y} = 292.308 - 0.214 x